OpenStudy (anonymous):

Find all values for z (complex number) where sin (z)=2. Find a branch of log(2z^2+1) that is analytic at i (imaginary number) adn compute the derivative at i

OpenStudy (experimentx):

try sin(z) = (e^z - e^-z)/2i = 2 try to find the value of e^z ... this will be quadratic.

OpenStudy (experimentx):

find e^z from there and then solve for z taking logs.

OpenStudy (anonymous):

\[sinz=\sin(a+ib)=cosasinib+cosibsina=sinxcoshy+isinhycosx=2\]

OpenStudy (anonymous):

\[*=isinh(b)cos(a)+\cosh(b)\sin(a)\]

OpenStudy (anonymous):

\[\sinh(b)\cos(a)=0\] \[\cosh(b)\sin(a)=2\]

OpenStudy (anonymous):

Where these multivalued functions intersect is your allowable values for a and b.

OpenStudy (anonymous):

Thank you for this